# Rates of convergence in long time asymptotics of an alignment model with symmetry breaking

### v2

@@ 3.3. Case $D<D_{*}$
$\left|\frac{du_{f}}{dt}\right|\leq N_{\beta}t^{-\frac{\beta}{2p}}.$ (31)
Integrating this inequality on $(t,+\infty)$, we obtain Proposition˜4.1 with $u_{\infty}\in\mathcal{S}$ because of Lemma˜2.2. ∎
@@
Let $u_{\infty}=\lim_{t\rightarrow\infty}u_{f}(t)$ which is well defined by Proposition˜4.1. Using (15), that is,
$\mathcal{H}[f(t,.)|G_{u_{\infty}}]=\frac{1}{D}\left(\mathcal{F}[f(t,.)]-\mathcal{F}_{*}+\frac{1}{2}|u_{f}-u_{*}|^{2}\right),$
@@
We know from the previous section that there exists an average speed limit $u_{\infty}\in\mathcal{S}$ for $u_{f}$ and that $f$ converges in relative entropy to $G_{u_{\infty}}$. It is natural to consider the linearized evolution operator around the steady state $G_{u_{\infty}}$. However, as it is mentioned in [11], the linearized operator around $G_{u_{\infty}}$ does not have a spectral gap if $u_{f}$ converges to $u_{\infty}$ tangentially to $\mathcal{S}$. We will avoid this problem by considering the evolution around $G_{*}:=G_{u_{*}}$ with $u_{*}$ defined by (30). Note that $G_{*}$ depends on $t$.
@@
First of all, let us define the quadratic forms associated to the free energy $\mathcal{F}$ and the Fisher information $\mathcal{I}$ as in [11]. For $u\in\mathcal{S}$ and
$g\in X_{u}:=\left\{g\in L^{2}(\mathbb{R}^{d},G_{u}dv)\text{ :}\int_{\mathbb{R}^{d}}g\,G_{u_{\,}}dv=0\right\},$
@@
  • $Q_{1,u}[g]:=\lim_{\epsilon\rightarrow 0}{\frac{2}{\epsilon^{2}}\big{(}\mathcal{F}[G_{u}(1+\epsilon g)]-\mathcal{F}_{*}\big{)}}=D\int_{\mathbb{R}^{d}}g^{2}G_{u}dv-D^{2}|v_{g}|^{2},$
  • $Q_{2,u}[g]:=\lim_{\epsilon\rightarrow 0}{\frac{1}{\epsilon^{2}}\mathcal{I}[G_{u}(1+\epsilon g)]}=D^{2}\int_{\mathbb{R}^{d}}|\nabla g-v_{g}|^{2}G_{u}dv,$
@@
$v_{g}=\frac{1}{D}\int_{\mathbb{R}^{d}}vgG_{u}dv.$ (32)
For $u\in\mathcal{S}$ fixed, we define on $X_{u}$ the bilinear form associated to the quadratic form $Q_{1,u}$:
$\langle g_{1},g_{2}\rangle_{u}=D\int_{\mathbb{R}^{d}}g_{1}g_{2}G_{u}dv-D^{2}v_{g_{1}}v_{g_{2}}.$ (33)
Since $G_{u}$ is a minimum of $\mathcal{F}$, the bilinear form $\langle\cdot,\cdot\rangle_{u}$ is nonnegative. We denote by
$\lVert g\rVert_{2,u}:=\sqrt{\int_{\mathbb{R}^{d}}g^{2}\,G_{u}\,dv}$ (34)
the classical weighted $L^{2}$ norm on $X_{u}$.
@@
Let $u\in\mathcal{S}$. Then, $f$ is a solution of (1) if and only if $g=(f-G_{u})G_{u}^{-1}$ solves
$\partial_{t}g=\mathcal{L}_{u}g+\mathcal{R}_{u}g$ (35)
@@
$\mathcal{L}_{u}g=D\Delta g+(\nabla\psi_{\alpha}+v-u)(v_{g}-\nabla g)$ (36)
@@
$\mathcal{R}_{u}g=-v_{g}\cdot[D\nabla g-(\nabla\psi_{\alpha}+v-u)g].$ (37)
Here, $\mathcal{L}_{u}$ is a linear operator defined on $X_{u}$ and $\mathcal{R}_{u}$ is the nonlinear remaining term. We recall that  $v_{g}$ is defined by (32) and $\psi_{\alpha}$ by (1).
@@
If $g_{\mathrm{ini}}\in X_{u}$. Then there exists a unique distributional solution of (35) in $X_{T}:=L^{2}(0,T;V)\cap C([0,T),H)\cap H^{1}(0,T;V^{\prime})$ with $H=X_{u}$ and $V=H^{1}(\mathbb{R}^{d},G_{u}\,dx)$.
An important observation is that the linear operator $\mathcal{L}_{u}$ is symmetric with respect to the scalar product $\langle\cdot,\cdot\rangle_{u}$. More precisely we have the following lemma which follows from several integration by parts.
@@
For all $u\in\mathcal{S}$ and $g,h\in H^{1}(\mathbb{R}^{d},G_{u}\,dv)$, we have,
$\langle\mathcal{L}_{u}g,h\rangle_{u}=-D^{2}\int_{\mathbb{R}^{d}}(\nabla g-v_{g})\cdot(\nabla h-v_{h})\,G_{u}\,dv\text{ and }\langle\mathcal{R}_{u}g,h\rangle_{u}=D^{2}\,v_{g}\cdot\int_{\mathbb{R}^{d}}g\,(\nabla h-v_{h})\,G_{u}\,dv.$
@@ 5.2. A coercivity result
Consider $u\in\mathcal{S}$. Assume that $g\in H^{1}(\mathbb{R}^{d},G_{u}\,dv)$ and $\int_{\mathbb{R}^{d}}g\,G_{u}\,dv=0$ then
$Q_{2,u}[g]\geq\beta(D)^{2}\frac{(v_{g}\cdot{u})^{2}}{|v_{g}|^{2}|u|^{2}}Q_{1,u}[g]$
where $\beta(D):=D\Lambda(D)\big{(}1-\kappa(D)\big{)}$ and
$\kappa(D):=\frac{1}{r(D)^{2}}\int_{\mathbb{R}^{d}}|(v-u)\cdot u|^{2}G_{u}dv\in(0,1)$
for an arbitrary $u\in\mathcal{S}$.
Lemma˜5.4 was proved in [11, Lemma 3.1] for $D>D_{*}$ and $u=0$. Here we adapt the proof for $D<D_{*}$.
@@
Let $u\in\mathcal{S}$ and $g\in L^{2}(\mathbb{R}^{d},G_{u}\,dv)$ such that $\int_{\mathbb{R}^{d}}g\,G_{u}\;dv=0$. Then
$Q_{1,u}[g]$ $\geq\eta(D)^{2}\frac{(v_{g}\cdot{u})^{2}}{|v_{g}|^{2}|u|^{2}}\int_{\mathbb{R}^{d}}g^{2}G_{u}dv,$
@@
If $w=\int_{\mathbb{R^{d}}}g\,v\,G_{u}\,dv=D\,v_{g}$ for some $u\in\mathcal{S}$, then
$|w|^{2}$ $=\int_{\mathbb{R}^{d}}{g(v-u)\cdot w\,G_{u}\,dv}\leq\lVert g\rVert_{2,u}\left(\int_{\mathbb{R}^{d}}|(v-u).w|^{2}G_{u}\right)^{\frac{1}{2}}.$
@@
$\frac{1}{D}\int_{\mathbb{R}^{d}}|v-u\cdot w|^{2}G_{u}dv=(\kappa(D)-1)(w\cdot e)^{2}+|w|^{2}\text{ where }e=u/r(D).$
@@
$|w|^{2}$ $\leq\lVert g\rVert_{2,u}\left[D\left((\kappa(D)-1)|w.e|^{2}+|w|^{2}\right)\right]^{\frac{1}{2}}$
$\leq\lVert g\rVert_{2,u}\left[D\left(1-(1-\kappa(D))\frac{|w\cdot e|^{2}}{|w|^{2}}\right)\right]^{\frac{1}{2}}|w|,$
@@
$\left|\int_{\mathbb{R^{d}}}{gvG_{u}dv}\right|^{2}\leq D\left(1-(1-\kappa(D))\frac{|w\cdot e|^{2}}{|w|^{2}}\right)\lVert g\rVert_{2,u}^{2},$
$Q_{1,u}[g]$ $=D\int_{\mathbb{R}^{d}}g^{2}G_{u}dv-\left|\int_{\mathbb{R}^{d}}gvG_{u}dv\right|^{2}\geq D(1-\kappa(D))\frac{|w\cdot e|^{2}}{|w|^{2}}\lVert g\rVert_{2,u}^{2}.$
@@
$W_{k}:=\int_{\mathbb{R}^{d}}|v|^{k}G_{u}dv$ (38)
for $u\in\mathcal{S}$ and observe that $W_{k}$ is independent of the chosen $u\in\mathcal{S}$.
@@
There exist positive constants $a(D)$ and $b(D)$ such that for all  $g\in H^{1}(\mathbb{R}^{d},G_{u}\,dv)$  satisfying $\int_{\mathbb{R}^{d}}gG_{u}\,dv=0$,
$a(D)Q_{2,u}[g]^{\frac{1}{2}}\leq\lVert\nabla g\rVert_{2,u}\leq b(D)\frac{|v_{g}|^{2}|u|^{2}}{(v_{g}\cdot u)^{2}}Q_{2,u}[g]^{\frac{1}{2}}.$
@@
$Q_{2,u}[g]$ $=D^{2}\lVert\nabla g-v_{g}\rVert_{2,u}^{2}$ (39)
$=D^{2}\lVert\nabla g\rVert_{2,u}^{2}-2D^{2}\,v_{g}\cdot\int_{\mathbb{R}^{d}}\nabla g\,G_{u}\,dv+D^{2}|v_{g}|^{2}$
$\geq D^{2}\lVert\nabla g\rVert_{2,u}^{2}-2DW_{2}^{1/2}\lVert g\rVert_{2,u}\lVert\nabla g\rVert_{2,u}$
@@
$P\left(\lVert\nabla g\rVert_{2,u}\right)\leq 0,$
@@
$P(X):=D^{2}X^{2}-2DW_{2}^{\frac{1}{2}}\lVert g\rVert_{2,u}X-Q_{2,u}[g].$
Using Lemma˜5.3 and Lemma˜5.4, we obtain
$\lVert\nabla g\rVert_{2,u}\leq\frac{2}{D}W_{2}^{\frac{1}{2}}\left(\lVert g\rVert_{2,u}+Q_{2,u}[g]^{\frac{1}{2}}\right)\leq b(D)\frac{|v_{g}|^{2}|u|^{2}}{(v_{g}\cdot u)^{2}}Q_{2,u}[g]^{\frac{1}{2}},$
@@
$b(D)=\frac{2}{D}W_{2}^{\frac{1}{2}}\left(\frac{1}{\eta(D)\beta(D)}+1\right).$
@@
$a(D)^{2}Q_{2,u}[g]\leq\lVert\nabla g\rVert_{2,u}^{2}$ (40)
@@
$\mathcal{A}_{\epsilon}:=\left\{g\in L^{2}(\mathbb{R}^{d},G_{u}dv)\text{ : }\ \int_{\mathbb{R}^{d}}g\,G_{u}\,dv=0\ and\ |v_{g}\cdot u|\geq\epsilon|v_{g}|\right\}.$
Lemma˜5.4 states that the norms $Q_{1,u}$ and $\lVert.\rVert_{2,u}$ are equivalent on $\mathcal{A}_{\epsilon}$. Lemma˜5.5 states that the norms associated to $Q_{2,u}$ and the canonical norm on $H^{1}(\mathbb{R}^{d},G_{u}\,dv)$ are equivalent on $\mathcal{A}_{\epsilon}$. Let $u\in\mathcal{S}$ and $g$ be a solution of (35). Using Lemma˜5.2, we obtain
$\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{u}$ $=\langle\partial_{t}g,g\rangle_{u}=\langle\mathcal{L}g+\mathcal{R}g,g\rangle_{u}=-Q_{2,u}[g]+D^{2}v_{g}\int_{\mathbb{R}^{d}}g(\nabla g-v_{g})G_{u}dv.$ (41)
@@
$\frac{1}{2}\frac{d}{dt}Q_{1,u}[g]=-\,Q_{2,u}[g].$
Since we proved $\lim_{t\rightarrow\infty}u_{f}(t)=u_{\infty}$ in Proposition˜4.1, a natural choice of $u$ is $u_{\infty}$. If $g$ belongs to some $\mathcal{A}_{\epsilon}$ with $u=u_{\infty}$ for all time $t\geq 0$, we are able to establish an exponential decay rate of $Q_{1,u_{\infty}}[g]$ thanks to Lemma˜5.4 and Grönwall’s lemma. Since $v_{g}=\frac{1}{D}(u_{f}-u_{\infty})$, this condition is equivalent to  $|(u_{f}-u_{\infty})\cdot u_{\infty}|\geq\epsilon|u_{f}-u_{\infty}|$. It is not satisfied if $u_{f}$ converges to $u_{\infty}$ tangentially to $\mathcal{S}$.
@@
where $u_{*}$ defined by (30) denotes the projection of $u_{f}$ on $\mathcal{S}$ and $G_{*}(t):=G_{u_{*}(t)}$. From now on, we take $u=u_{*}$. According to (32), $v_{g}=\frac{1}{D}\int_{\mathbb{R}^{d}}g\,v\,G_{*}\,dv.$
@@
Under the assumptions of Theorem˜1.2, if $f$ solves (1) and $g$ and $u_{*}$ are defined by (42) and (30), then, $g(t)$ is a function in $H^{1}(\mathbb{R}^{d},G_{*}(t)dv)$ for all $t>0$ and
$\partial_{t}g=\mathcal{L}_{*}g+\mathcal{R}_{*}g-\frac{v\cdot u_{*}^{\prime}}{D}(1+g)$ (43)
where $\mathcal{L}_{*}:=\mathcal{L}_{u_{*}}$ and $\mathcal{R}_{*}:=\mathcal{R}_{u_{*}}$ are defined by (36) and (37). Moreover there holds $u_{*}^{\prime}\cdot v_{g}=0$ where $u_{*}^{\prime}:=\frac{\mathrm{d}u_{*}}{\mathrm{d}t}$
@@
Let $t_{0}>0$ and $u_{0}:=u_{*}(t_{0})$. Let $f(t)=G_{u_{0}}(1+g^{0}(t))$ for $t\in\mathbb{R}$. By assumption of Theorem˜1.2, $f_{\mathrm{ini}}\in\cap_{u\in\mathcal{S}}L^{2}(\mathbb{R}^{d},G_{u}^{-1}dv)\subset L^{2}(\mathbb{R}^{d},G_{u_{0}}^{-1}dv)$. Hence, $g^{0}_{\mathrm{ini}}:=g^{0}(0)\in L^{2}(\mathbb{R}^{d},G_{u_{0}}dv).$ Thanks to Proposition˜5.1 we get $g^{0}\in X_{T}$ hence,
$g(t_{0})=g^{0}(t_{0})\in H^{1}(\mathbb{R}^{d},G_{u_{0}}dv).$

### v3

@@ 3.3. Case $D\geq D_{*}$
$\left|\frac{d\mathbf{u}_{f}}{dt}\right|\leq N_{\beta}t^{-\frac{\beta}{2p}}.$ (32)
Integrating this inequality on $(t,+\infty)$, we obtain Proposition˜4.1 with $\mathbf{u}_{\infty}\in\mathcal{S}$ because of Lemma˜2.2. ∎
@@
Let $\mathbf{u}_{\infty}=\lim_{t\rightarrow\infty}\mathbf{u}_{f}(t)$ which is well defined by Proposition˜4.1. Using (15), that is,
$\mathcal{H}[f(t,.)|G_{\mathbf{u}_{\infty}}]=\frac{1}{D}\left(\mathcal{F}[f(t,.)]-\mathcal{F}_{*}+\frac{1}{2}|\mathbf{u}_{f}-\mathbf{u}_{*}|^{2}\right),$
@@
We know from the previous section that there exists an average speed limit $\mathbf{u}_{\infty}\in\mathcal{S}$ for $\mathbf{u}_{f}$ and that $f$ converges in relative entropy to $G_{\mathbf{u}_{\infty}}$. It is natural to consider the linearized evolution operator around the steady state $G_{\mathbf{u}_{\infty}}$. However, as it is mentioned in [11], the linearized operator around $G_{\mathbf{u}_{\infty}}$ does not have a spectral gap if $\mathbf{u}_{f}$ converges to $\mathbf{u}_{\infty}$ tangentially to $\mathcal{S}$. We will avoid this problem by considering the evolution around $G_{*}:=G_{\mathbf{u}_{*}}$ with $\mathbf{u}_{*}$ defined by (31). Note that $G_{*}$ depends on $t$.
@@
First of all, let us define the quadratic forms associated to the free energy $\mathcal{F}$ and the Fisher information $\mathcal{I}$ as in [11]. For $\mathbf{u}\in\mathcal{S}$ and
$g\in X_{\mathbf{u}}:=\left\{g\in L^{2}(\mathbb{R}^{d},G_{\mathbf{u}}dv)\text{ :}\int_{\mathbb{R}^{d}}g\,G_{\mathbf{u}_{\,}}dv=0\right\},$
@@
  • $Q_{1,\mathbf{u}}[g]:=\lim_{\epsilon\rightarrow 0}{\frac{2}{\epsilon^{2}}\big(\mathcal{F}[G_{\mathbf{u}}(1+\epsilon g)]-\mathcal{F}_{*}\big)}=D\int_{\mathbb{R}^{d}}g^{2}G_{\mathbf{u}}dv-D^{2}|\mathbf{v}_{g}|^{2},$
  • $Q_{2,\mathbf{u}}[g]:=\lim_{\epsilon\rightarrow 0}{\frac{1}{\epsilon^{2}}\mathcal{I}[G_{\mathbf{u}}(1+\epsilon g)]}=D^{2}\int_{\mathbb{R}^{d}}|\nabla g-\mathbf{v}_{g}|^{2}G_{\mathbf{u}}dv,$
@@
$\mathbf{v}_{g}=\frac{1}{D}\int_{\mathbb{R}^{d}}vgG_{\mathbf{u}}dv.$ (33)
For $\mathbf{u}\in\mathcal{S}$ fixed, we define on $X_{\mathbf{u}}$ the bilinear form associated to the quadratic form $Q_{1,\mathbf{u}}$:
$\langle g_{1},g_{2}\rangle_{\mathbf{u}}=D\int_{\mathbb{R}^{d}}g_{1}g_{2}G_{\mathbf{u}}dv-D^{2}v_{g_{1}}v_{g_{2}}.$ (34)
Since $G_{\mathbf{u}}$ is a minimum of $\mathcal{F}$, the bilinear form $\langle\cdot,\cdot\rangle_{\mathbf{u}}$ is nonnegative. We denote by
$\lVert g\rVert_{2,\mathbf{u}}:=\sqrt{\int_{\mathbb{R}^{d}}g^{2}\,G_{\mathbf{u}}\,dv}$ (35)
the classical weighted $L^{2}$ norm on $X_{\mathbf{u}}$. The following lemma is proved in [11, Section 5.1] for $\mathbf{u}=0$. By adapting the computations, we get the following result for any $\mathbf{u}\in\mathbb{R}^{d}.$
@@
Let $\mathbf{u}\in\mathcal{S}$. Then, $f$ is a solution of (1) if and only if $g=(f-G_{\mathbf{u}})G_{\mathbf{u}}^{-1}$ solves
$\partial_{t}g=\mathcal{L}_{\mathbf{u}}g+\mathcal{R}_{\mathbf{u}}g$ (36)
@@
$\mathcal{L}_{\mathbf{u}}g=D\Delta g+(\nabla\psi_{\alpha}+v-\mathbf{u})(\mathbf{v}_{g}-\nabla g)$ (37)
@@
$\mathcal{R}_{\mathbf{u}}g=-\mathbf{v}_{g}\cdot[D\nabla g-(\nabla\psi_{\alpha}+v-\mathbf{u})g].$ (38)
Here, $\mathcal{L}_{\mathbf{u}}$ is a linear operator defined on $X_{\mathbf{u}}$ and $\mathcal{R}_{\mathbf{u}}$ is the nonlinear remaining term. We recall that $\mathbf{v}_{g}$ is defined by (33) and $\psi_{\alpha}$ by (1).
@@
If $g_{\mathrm{in}}\in X_{\mathbf{u}}$. Then there exists a unique distributional solution of (36) in $X_{T}:=L^{2}(0,T;V)\cap C([0,T),H)\cap H^{1}(0,T;V^{\prime})$ with $H=X_{\mathbf{u}}$ and $V=H^{1}(\mathbb{R}^{d},G_{\mathbf{u}}\,dx)$.
An important observation is that the linear operator $\mathcal{L}_{\mathbf{u}}$ is symmetric with respect to the scalar product $\langle\cdot,\cdot\rangle_{\mathbf{u}}$. More precisely we have the following lemma which follows from several integration by parts and is detailed in [11, Lemma 5.1].
@@
For all $\mathbf{u}\in\mathcal{S}$ and $g,h\in H^{1}(\mathbb{R}^{d},G_{\mathbf{u}}\,dv)$, we have,
$\langle\mathcal{L}_{\mathbf{u}}g,h\rangle_{\mathbf{u}}=-D^{2}\int_{\mathbb{R}^{d}}(\nabla g-\mathbf{v}_{g})\cdot(\nabla h-v_{h})\,G_{\mathbf{u}}\,dv\text{ and }\langle\mathcal{R}_{\mathbf{u}}g,h\rangle_{\mathbf{u}}=D^{2}\,\mathbf{v}_{g}\cdot\int_{\mathbb{R}^{d}}g\,(\nabla h-v_{h})\,G_{\mathbf{u}}\,dv.$
@@ 5.2. A coercivity result
Consider $\mathbf{u}\in\mathcal{S}$. Assume that $g\in H^{1}(\mathbb{R}^{d},G_{\mathbf{u}}\,dv)$ and $\int_{\mathbb{R}^{d}}g\,G_{\mathbf{u}}\,dv=0$ then
$Q_{2,\mathbf{u}}[g]\geq\beta(D)^{2}\frac{(\mathbf{v}_{g}\cdot{\mathbf{u}})^{2}}{|\mathbf{v}_{g}|^{2}|\mathbf{u}|^{2}}Q_{1,\mathbf{u}}[g]$
where $\beta(D)^{2}:=D\Lambda(D)\big(1-\kappa(D)\big)$ and
$\kappa(D):=\frac{1}{r(D)^{2}}\int_{\mathbb{R}^{d}}|(v-\mathbf{u})\cdot\mathbf{u}|^{2}G_{\mathbf{u}}dv\in(0,1)$
for an arbitrary $\mathbf{u}\in\mathcal{S}$.
Lemma˜5.4 was proved in [11, Lemma 3.1] for $D>D_{*}$ and $\mathbf{u}=0$. Here we adapt the proof for $D<D_{*}$.
@@
Let $\mathbf{u}\in\mathcal{S}$ and $g\in L^{2}(\mathbb{R}^{d},G_{\mathbf{u}}\,dv)$ such that $\int_{\mathbb{R}^{d}}g\,G_{\mathbf{u}}\;dv=0$. Then
$Q_{1,\mathbf{u}}[g]$ $\geq\eta(D)^{2}\frac{(\mathbf{v}_{g}\cdot{\mathbf{u}})^{2}}{|\mathbf{v}_{g}|^{2}|\mathbf{u}|^{2}}\int_{\mathbb{R}^{d}}g^{2}G_{\mathbf{u}}dv,$
@@
If $w=\int_{\mathbb{R^{d}}}g\,v\,G_{\mathbf{u}}\,dv=D\,\mathbf{v}_{g}$ for some $\mathbf{u}\in\mathcal{S}$, then
$|w|^{2}$ $=\int_{\mathbb{R}^{d}}{g(v-\mathbf{u})\cdot w\,G_{\mathbf{u}}\,dv}\leq\lVert g\rVert_{2,\mathbf{u}}\left(\int_{\mathbb{R}^{d}}|(v-\mathbf{u}).w|^{2}G_{\mathbf{u}}\,dv\right)^{\frac{1}{2}}.$
@@
$\frac{1}{D}\int_{\mathbb{R}^{d}}|v-\mathbf{u}\cdot w|^{2}G_{\mathbf{u}}dv=(\kappa(D)-1)(w\cdot e)^{2}+|w|^{2}\text{ where }e=\mathbf{u}/r(D).$
@@
$|w|^{2}$ $\leq\lVert g\rVert_{2,\mathbf{u}}\left[D\left((\kappa(D)-1)|w.e|^{2}+|w|^{2}\right)\right]^{\frac{1}{2}}$
$\leq\lVert g\rVert_{2,\mathbf{u}}\left[D\left(1-(1-\kappa(D))\frac{|w\cdot e|^{2}}{|w|^{2}}\right)\right]^{\frac{1}{2}}|w|,$
@@
$\left|\int_{\mathbb{R^{d}}}{gvG_{\mathbf{u}}dv}\right|^{2}\leq D\left(1-(1-\kappa(D))\frac{|w\cdot e|^{2}}{|w|^{2}}\right)\lVert g\rVert_{2,\mathbf{u}}^{2},$
$Q_{1,\mathbf{u}}[g]$ $=D\int_{\mathbb{R}^{d}}g^{2}G_{\mathbf{u}}dv-\left|\int_{\mathbb{R}^{d}}gvG_{\mathbf{u}}dv\right|^{2}\geq D(1-\kappa(D))\frac{|w\cdot e|^{2}}{|w|^{2}}\lVert g\rVert_{2,\mathbf{u}}^{2}.$
@@
$W_{k}:=\int_{\mathbb{R}^{d}}|v|^{k}G_{\mathbf{u}}dv$ (39)
for $\mathbf{u}\in\mathcal{S}$ and observe that $W_{k}$ is independent of the chosen $\mathbf{u}\in\mathcal{S}$.
@@
There exist positive constants $a(D)$ and $b(D)$ such that for all $g\in H^{1}(\mathbb{R}^{d},G_{\mathbf{u}}\,dv)$ satisfying $\int_{\mathbb{R}^{d}}gG_{\mathbf{u}}\,dv=0$,
$a(D)Q_{2,\mathbf{u}}[g]^{\frac{1}{2}}\leq\lVert\nabla g\rVert_{2,\mathbf{u}}\leq b(D)\frac{|\mathbf{v}_{g}|^{2}|\mathbf{u}|^{2}}{(\mathbf{v}_{g}\cdot\mathbf{u})^{2}}Q_{2,\mathbf{u}}[g]^{\frac{1}{2}}.$
@@
$Q_{2,\mathbf{u}}[g]$ $=D^{2}\lVert\nabla g-\mathbf{v}_{g}\rVert_{2,\mathbf{u}}^{2}$ (40)
$=D^{2}\lVert\nabla g\rVert_{2,\mathbf{u}}^{2}-2D^{2}\,\mathbf{v}_{g}\cdot\int_{\mathbb{R}^{d}}\nabla g\,G_{\mathbf{u}}\,dv+D^{2}|\mathbf{v}_{g}|^{2}$
$\geq D^{2}\lVert\nabla g\rVert_{2,\mathbf{u}}^{2}-2DW_{2}^{1/2}\lVert g\rVert_{2,\mathbf{u}}\lVert\nabla g\rVert_{2,\mathbf{u}}$
@@
$P\left(\lVert\nabla g\rVert_{2,\mathbf{u}}\right)\leq 0,$
@@
$P(X):=D^{2}X^{2}-2DW_{2}^{\frac{1}{2}}\lVert g\rVert_{2,\mathbf{u}}X-Q_{2,\mathbf{u}}[g].$
Using that for $a>0,b>0,c>0,$
$ax^{2}-bx-c\leq 0\Rightarrow x\leq\frac{b+\sqrt{b^{2}+4ac}}{2a}\leq\frac{b}{a}+\sqrt{\frac{c}{a}}$
with Lemma˜5.3 and Lemma˜5.4, we obtain
$\lVert\nabla g\rVert_{2,\mathbf{u}}\leq\frac{1}{D}\left(2W_{2}^{\frac{1}{2}}\lVert g\rVert_{2,\mathbf{u}}+Q_{2,\mathbf{u}}[g]^{\frac{1}{2}}\right)\leq b(D)\frac{|\mathbf{v}_{g}|^{2}|\mathbf{u}|^{2}}{(\mathbf{v}_{g}\cdot\mathbf{u})^{2}}Q_{2,\mathbf{u}}[g]^{\frac{1}{2}},$
@@
$b(D)=\frac{1}{D}\left(\frac{2W_{2}^{\frac{1}{2}}}{\eta(D)\beta(D)}+1\right).$
@@
$a(D)^{2}Q_{2,\mathbf{u}}[g]\leq\lVert\nabla g\rVert_{2,\mathbf{u}}^{2}$ (41)
@@
$\mathcal{A}_{\epsilon}:=\left\{g\in L^{2}(\mathbb{R}^{d},G_{\mathbf{u}}dv)\text{ : }\ \int_{\mathbb{R}^{d}}g\,G_{\mathbf{u}}\,dv=0\ and\ |\mathbf{v}_{g}\cdot\mathbf{u}|\geq\epsilon|\mathbf{v}_{g}|\right\}.$
Lemma˜5.4 states that the norms $Q_{1,\mathbf{u}}$ and $\lVert.\rVert_{2,\mathbf{u}}$ are equivalent on $\mathcal{A}_{\epsilon}$. Lemma˜5.5 states that the norms associated to $Q_{2,\mathbf{u}}$ and the canonical norm on $H^{1}(\mathbb{R}^{d},G_{\mathbf{u}}\,dv)$ are equivalent on $\mathcal{A}_{\epsilon}$. Let $\mathbf{u}\in\mathcal{S}$ and $g$ be a solution of (36). Using Lemma˜5.2, we obtain
$\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{\mathbf{u}}$ $=\langle\partial_{t}g,g\rangle_{\mathbf{u}}=\langle\mathcal{L}g+\mathcal{R}g,g\rangle_{\mathbf{u}}=-Q_{2,\mathbf{u}}[g]+D^{2}\mathbf{v}_{g}\int_{\mathbb{R}^{d}}g(\nabla g-\mathbf{v}_{g})G_{\mathbf{u}}dv.$ (42)
@@
$\frac{1}{2}\frac{d}{dt}Q_{1,\mathbf{u}}[g]=-\,Q_{2,\mathbf{u}}[g].$
Since we proved $\lim_{t\rightarrow\infty}\mathbf{u}_{f}(t)=\mathbf{u}_{\infty}$ in Proposition˜4.1, a natural choice of $\mathbf{u}$ is $\mathbf{u}_{\infty}$. If $g$ belongs to some $\mathcal{A}_{\epsilon}$ with $\mathbf{u}=\mathbf{u}_{\infty}$ for all time $t\geq 0$, we are able to establish an exponential decay rate of $Q_{1,\mathbf{u}_{\infty}}[g]$ thanks to Lemma˜5.4 and Grönwall’s lemma. Since $\mathbf{v}_{g}=\frac{1}{D}(\mathbf{u}_{f}-\mathbf{u}_{\infty})$, this condition is equivalent to $|(\mathbf{u}_{f}-\mathbf{u}_{\infty})\cdot\mathbf{u}_{\infty}|\geq\epsilon|\mathbf{u}_{f}-\mathbf{u}_{\infty}|$. It is not satisfied if $\mathbf{u}_{f}$ converges to $\mathbf{u}_{\infty}$ tangentially to $\mathcal{S}$.
@@
where $\mathbf{u}_{*}$ defined by (31) denotes the projection of $\mathbf{u}_{f}$ on $\mathcal{S}$ and $G_{*}(t):=G_{\mathbf{u}_{*}(t)}$. From now on, we take $\mathbf{u}=\mathbf{u}_{*}$. According to (33), $\mathbf{v}_{g}=\frac{1}{D}\int_{\mathbb{R}^{d}}g\,v\,G_{*}\,dv.$
@@
Under the assumptions of Theorem˜1.2, if $f$ solves (1) and $g$ and $\mathbf{u}_{*}$ are defined by (43) and (31), then, $g(t)$ is a function in $H^{1}(\mathbb{R}^{d},G_{*}(t)dv)$ for all $t>0$ and
$\partial_{t}g=\mathcal{L}_{*}g+\mathcal{R}_{*}g-\frac{v\cdot\mathbf{u}_{*}^{\prime}}{D}(1+g)$ (44)
where $\mathcal{L}_{*}:=\mathcal{L}_{\mathbf{u}_{*}}$ and $\mathcal{R}_{*}:=\mathcal{R}_{\mathbf{u}_{*}}$ are defined by (37) and (38). Moreover there holds $\mathbf{u}_{*}^{\prime}\cdot\mathbf{v}_{g}=0$ where $\mathbf{u}_{*}^{\prime}:=\frac{\mathrm{d}\mathbf{u}_{*}}{\mathrm{d}t}$
@@
Let $t_{0}>0$ and $\mathbf{u}_{0}:=\mathbf{u}_{*}(t_{0})$. Let $f(t)=G_{\mathbf{u}_{0}}(1+g^{0}(t))$ for $t\in\mathbb{R}$. By assumption of Theorem˜1.2, $f_{\mathrm{in}}\in\cap_{\mathbf{u}\in\mathcal{S}}L^{2}(\mathbb{R}^{d},G_{\mathbf{u}}^{-1}dv)\subset L^{2}(\mathbb{R}^{d},G_{\mathbf{u}_{0}}^{-1}dv)$. Hence, $g^{0}_{\mathrm{ini}}:=g^{0}(0)\in L^{2}(\mathbb{R}^{d},G_{\mathbf{u}_{0}}dv).$ Thanks to Proposition˜5.1 we get $g^{0}\in X_{T}$ hence,
$g(t_{0})=g^{0}(t_{0})\in H^{1}(\mathbb{R}^{d},G_{\mathbf{u}_{0}}dv).$